Because the location of the curved super optic has changed inside the cavity from its initial position, we must change the mode matching for the 1550nm light in the filter cavity. Previously, the waist was calculated to be at the third mirror the 1550nm light encounters, whereas now the waist is on the input coupler. The new MM solution is approximately f1 = .250 @ .902m and f2 = 1.5 @ 1.546m, where the distances are counting from the 1550nm fiber collimator. I placed new lenses according to the solution and measured the beam to be slightly too big at the measure point. This means L1 was too far back, it was moved forward <1 inch to match the predicted beam size. We can see the first pass beam on the SWIR camera but alignment will need to be improved before we see any coresonance.